Evaluate f(x)=4x+3 for f(-5)
Answer by Theo(13342)

f(-5) measn to evaluate the function when the value of x is equal to -5.

f(x) = 4x + 3

f(-5) = 4*-5 + 3 = -20 + 3 = -17